OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches two of quantity x squared minus four divided by quantity x minus two.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}\frac{ x ^{2}-4 }{ x-2 }\]

OpenStudy (anonymous):

@ganeshie8, @amistre64

OpenStudy (anonymous):

is it 0 or does it not exist?

OpenStudy (anonymous):

it exists and not 0

OpenStudy (anonymous):

factorize the numerator and then proceed.

OpenStudy (anonymous):

could you show me how to do it?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}\frac{ \left( x+2 \right)\left( x-2 \right) }{ \left( x-2 \right) }=\lim_{x \rightarrow 2}\left( x+2 \right)\] =?

OpenStudy (anonymous):

is it then =4?

OpenStudy (anonymous):

you are correct.